OpenStudy (anonymous):

How do I factor 12b^2 + 25b - 7 ? I've completely forgotten what I'm supposed to do and my book isn't particularly helping.

OpenStudy (whpalmer4):

multiply 12 and -7, what do you get?

OpenStudy (anonymous):

-84 ?

OpenStudy (whpalmer4):

yes. now, what are two factors of -84 that add to 25?

OpenStudy (anonymous):

-3 and 28 ?

RhondaSommer (rhondasommer):

Multiply 12 by -7, which is -84. Which two numbers add up to 25 and multiply to -84? 28 and -3 Rewrite 25b as the sum of 28b and −3b

OpenStudy (whpalmer4):

yes. this means we can split the 25b term into -3b and 28 b: \[12b^2+28b-3b-7\] now we need to take a pair of those terms so that we can factor it. can't factor \(3b-7\) so we will do \(28b-7\) which we can factor as \[7(4b-1)\] how about the other two terms?

RhondaSommer (rhondasommer):

no because you are suppposed to divide the largest number possible being 4b

OpenStudy (whpalmer4):

\[12b^2-3b=3b(4b-1)\] notice that the bit in the ( ) is identical to that in our other term? \[12b^2+28b-3b-7= (12b^2-3b) + (28b-7) = 3b(4b-1)+7(4b-1)\]\[=(3b+7)(4b-1)\]

OpenStudy (anonymous):

So from 12b^2+28b−3b−7 I can factor by grouping and get 3b(4b-1) + 7(4b-1) which is the same as (3b+7)(4b-1) right? And that'd be the answer?

OpenStudy (whpalmer4):

checking our work: \[(3b+7)(4b-1)=(3b)(4b)+(-1)(3b)+7(4b)+7(-1)=12b^2+28b-3b-7\]\[=12b^+25b-7\checkmark\]

OpenStudy (whpalmer4):

yes, mangled the formatting a bit on my last post, but your answer is correct

OpenStudy (anonymous):

Okay ! Thank you a ton, I forgot what to do and I was super confused. I think i get it now though. Again, thanks !

OpenStudy (whpalmer4):

I always have to figure it out again when I haven't done it for a while!